# Emotion Granularity from Text: An Aggregate-Level Indicator of Mental Health

### v1

@@ Appendix C Emotion Granularity: Various Window Sizes
    We report the results of the statistical analyses of emotion granularity when emotion arcs were generated using two other window sizes: 100 (Table 3) and 500 (Table 4). We find that largely the results do not change, however there are some differences in the scenario when the dataset was smaller (e.g., eRisk dataset or MHC such as MDD). In such cases, when the window size is increased, it is possible that several emotional experiences occurred, resulting in a weaker signal of emotion granularity.
@@
    Appendix D Emotion Granularity: Emotion Pairs
@@
    anger–anticipation lower lower lower lower lower lower
    anger–disgust lower lower lower lower lower lower lower lower
    anger–fear lower lower lower lower lower
    anger–joy lower higher higher higher higher higher
    anger–sadness lower lower lower lower lower lower lower lower
    anger–surprise lower higher lower higher lower
    anger–trust lower lower lower lower lower
@@
    anticipation–fear lower lower lower lower lower lower lower lower
    anticipation–joy lower lower higher
    anticipation–sadness lower lower lower lower lower lower lower lower
    anticipation–surprise higher lower lower lower lower lower lower lower
    anticipation–trust lower lower lower lower lower
@@
    disgust–joy lower higher higher higher
    disgust–sadness lower lower lower lower lower lower
    disgust–surprise lower lower lower lower higher lower
    disgust–trust lower lower lower lower lower
    fear–joy lower
    fear–sadness lower lower lower lower lower lower lower lower
    fear–surprise lower higher higher higher
    fear–trust lower lower lower lower lower lower
    joy–sadness lower higher
    joy–surprise higher lower lower
@@
    sadness–surprise lower higher higher higher higher lower
    sadness–trust lower lower lower lower lower
    surprise–trust lower lower lower lower lower lower lower
    Table 5: Emotion Granularity - Utterance-Level: The difference in emotion granularity between each emotion pair, for each MHC group and the control in the Twitter-STMHD dataset. A significant difference is indicated by the word ‘lower’ or ‘higher’, indicating the direction of the difference.
    Appendix E Term Specificity Results
@@
    Appendix F Emotion Correlations
@@
    Table 8 shows the group-averaged Spearman correlations for emotion pairs in the positive, negative, mixed valence groups, and the within-group and cross-group averages, for the MHCs and Control groups in both datasets.
    emo2 anticipation disgust fear joy sadness surprise trust
    emo1
    anger 0.020 0.593 0.536 -0.015 0.447 0.113 0.003
    anticipation 0.009 0.040 0.432 0.022 0.251 0.404
    disgust 0.429 -0.014 0.434 0.061 0.001
    fear -0.010 0.567 0.138 0.004
    joy -0.009 0.221 0.467
    sadness 0.100 -0.001
    surprise 0.145
    Table 7: Emotion–Emotion Spearman Correlation: Spearman correlation values between pairs of utterance-level emotion arcs for the control users in the Twitter-STMHD dataset.
@@
        Control 0.467 0.501 0.251 0.083 0.465
        ADHD 0.495 0.524 0.243 0.096 0.486
        Anxiety 0.505 0.522 0.274 0.089 0.489
        Bipolar 0.491 0.524 0.273 0.084 0.488
        Depression 0.505 0.520 0.272 0.087 0.487
        MDD 0.508 0.518 0.267 0.085 0.486
        OCD 0.497 0.523 0.260 0.090 0.487
        PPD 0.473 0.511 0.269 0.081 0.476
        PTSD 0.485 0.521 0.259 0.090 0.484
@@
       Control 0.475 0.537 0.271 0.168 0.496
       Depression 0.521 0.568 0.319 0.184 0.531
    Table 8: Emotion Groups Spearman Correlation: Spearman correlation values between utterance-level emotion arcs of emotion pairs in the positive, negative, and variable valence groups, and within-group (overall), and cross-group averages, for the Twitter-STMHD and Reddit eRisk datasets. Emotion granularity is defined as the negative of these correlations (i.e, higher correlations imply a lower granularity).

### v2

@@ F.1 Various Window Sizes
    We find that largely the results do not change, however there are some differences in the scenario when the dataset was smaller (e.g., eRisk dataset or MHC such as MDD). In such cases, when the window size is increased, it is possible that several emotional experiences occurred, resulting in a weaker signal of emotion granularity.
@@
    Appendix G Emotion Granularity: Emotion Pairs
@@
    anger–anticipation lower lower lower lower lower
    anger–disgust lower lower lower lower lower lower lower
    anger–fear lower lower lower lower lower lower lower
    anger–joy lower lower higher lower lower
    anger–sadness lower lower lower lower lower lower lower
    anger–surprise lower lower lower lower lower
    anger–trust lower lower lower lower lower lower lower
@@
    anticipation–fear lower lower lower lower lower lower lower
    anticipation–joy lower lower lower lower lower lower lower lower
    anticipation–sadness lower lower lower lower lower lower
    anticipation–surprise lower lower lower lower lower lower
    anticipation–trust lower lower lower lower lower lower lower lower
@@
    disgust–joy lower lower lower lower lower
    disgust–sadness lower lower lower lower lower lower lower
    disgust–surprise lower lower lower lower lower lower
    disgust–trust lower lower lower lower lower lower lower lower
    fear–joy lower lower higher lower lower lower
    fear–sadness lower lower lower lower lower lower lower
    fear–surprise lower lower lower lower lower lower
    fear–trust lower lower lower lower lower lower lower lower
    joy–sadness lower lower lower lower lower
    joy–surprise lower lower lower lower lower
@@
    sadness–surprise lower lower lower lower lower lower lower
    sadness–trust lower lower lower lower lower lower lower lower
    surprise–trust lower lower lower lower lower lower
    Table 10: Emotion Granularity - emotion Pairs: The difference in emotion granularity between each emotion pair, for each MHC group and the control in the Twitter-STMHD dataset. A significant difference is indicated by the word ‘lower’ or ‘higher’, indicating the direction of the difference.
    Appendix H Term Specificity Results
@@
    Appendix I Emotion Correlations
    Table 12 shows the group-averaged Spearman correlations for emotion pairs in the positive, negative, mixed valence groups, and the within-group and cross-group averages, for the Control groups, and the delta from these values for each MHC in both datasets.
@@ Appendix I Emotion Correlations
        Control 0.027 0.023 0.012 0.006 0.022
        ADHD -0.012* -0.008* -0.005* -$0$010*
        Anxiety -0.012* -0.008* -0.003* -0.008* -0.010*
        Bipolar -0.004* -0.008* -0.004* -0.002* -0.006*
        MDD -0.011* -0.002 -0.001 -0.005* -0.005*
        OCD -0.013* -0.009* -0.006* -0.009* -0.009*
        PPD -0.005 -0.009* -0.005 -0.004 -0.003
        PTSD -0.013* -0.014* -0.006* -0.008* -0.013*
        Depression -0.011* -0.005* -0.003* -0.005* -0.008*
@@
       Control 0.114 0.117 0.090 0.094 0.112
       Depression -0.016* -0.021* -0.012 -0.022* -0.017*
    Table 12: Emotion Granularity - Spearman correlations: Spearman correlation values between utterance-level emotion arcs for the Control group, and the delta for each MHC when compared to the Control group. Emotion granularity is defined as the negative of these correlations (i.e, higher correlations imply a lower granularity). Hyperparameters are the same as in Table 2.
    anger anticipation disgust fear joy sadness surprise trust
    anger -0.003 0.020 0.027 -0.010 0.024 0.009 0.007
    anticipation -0.003 0.007 0.021 0.004 0.012 0.027
    disgust 0.023 -0.010 0.021 0.006 0.003
    fear -0.003 0.021 0.012 0.013
    joy -0.000 0.008 0.027
    sadness 0.012 0.008
    surprise 0.023
    Table 13: Emotion–Emotion Spearman correlations: Spearman correlation values between pairs of utterance-level emotion arcs for the all users in the control group of the Twitter-STMHD dataset. Hyperparameters are the same as in Table 2.